OpenStudy (anonymous):

The volume of a gas V held at a constant temperature in a closed container varies inversely with its pressure P. If the volume of a gas is 1500 cubic centimeters when the pressure is 450 mm Hg, find the volume when the pressure is 500 mm Hg.

OpenStudy (mathmate):

Use Boyle's law \[PV=constant\] or, equivalently \[P_1V_1 = P_2V_2\]

OpenStudy (abhisar):

\(\boxed{\boxed{\boxed{\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}}}}\)

OpenStudy (abhisar):

as @mathmate said use the equation P1V1 = P2V2, here P1 = 450 mmHg and V1 = 1500 ml & P2 = 500 mmHg

OpenStudy (abhisar):

*1cc = 1ml

OpenStudy (harsha19111999):

p1v1=p2v2 =>v2=p1v1/p2 =450*1500/500 = 450*3= 1350cc